OpenStudy (anonymous):

Estimate the area under the graph of y=x^3+3x for x between 0 and 3 . Use a partition that consists of 6 equal subintervals of [0,3] and use the right endpoint of each subinterval as a sample point.

OpenStudy (campbell_st):

create a table of function values  use either the Trapezoidal Rule, which finds the area of a trapeziod \[A = h/2 {(f(a) - f(b))}\] h = 1/2 f(a) = f(0) f(b) = f(1) this needs to be done 6 times... find the area of 6 trapezoids and add the answers to approximate the area under the curve. Alternative Simpsons Rule Simple form A \[A = h/6 {f(a) = 4\times f((a+b)/2) + f(b)} \] the more complex form if A = h/3( 1st + last + 2 x odds + 4 x evens} Trapezoidal rule is probably easier